The reaction 2A+3B=C is first order with respect to A and B. When the initial concentrations are A=1.6X10^-2 M and B= 2.4X10^-3 M. The rate is 4.1 X10^-4 M/s. Calculate the rate constant of the reaction.
1
Expert's answer
2017-04-20T06:49:42-0400
Solution:
r = k[A][B]
k = r/([A][B])= 4.1*10-4/(1.6*10-2*2.4*10-3)= 10.677 (M-1s-1)
Answer: the rate constant of the reaction is k = 10.677 M-1s-1